# Implementation of Leaking Quantum Walks on a Photonic Processor

### v1

@@ Abstract
Quantum walks represent pillars of quantum dynamics and information processing. They provide a powerful framework for simulating quantum transport, designing search algorithms, and achieving universal quantum computation. Several physical platforms have been employed to implement QWs, such as trapped atoms, trapped ions, nuclear magnetic resonance systems and photonic quantum systems either in bulk optics or waveguide structures and fiber-loop networks. Here we focus on the most promising approach, that is photonic integrated circuits.
We will review how the employment of this versatile experimental platform has allowed to explore several phenomena related to QW-based protocols, e.g. the evolution in presence of different kinds of noise. In this landscape, to the best of our knowledge, few examples report on the introduction of absorbing centers and their effects on the coherence of the dynamics. Here we present and discuss the results related to absorbing boundaries in QWs obtained through theoretical simulations and experiments conducted with the universal photonic quantum processors realized by Quix Quantum.
@@ 1 Introduction
In real photonic quantum systems, particle losses are unavoidable due to the non-ideal efficiencies of the network components and measurements apparatuses, often representing a major limitation for scaling up quantum applications. Nonetheless, the inclusion of absorbing sites within a QW remains a scenario that has received relatively little attention and has not yet been fully explored. In this regard, a limited number of numerical studies [7, 40, 50, 27, 4] and a few experimental implementations [30, 31] have been reported, showing how some recursive behavior is observed in the presence of specific absorbing sites. However, many aspects of this problem remain largely unexplored, even though the introduction of controlled losses appears to be a promising approach for probing the dynamics of open quantum systems. In this work, we address this gap by presenting both numerical simulations and experimental results for a QW in the presence of an absorbing boundary. Specifically, we investigate a walk on a finite lattice featuring a partially absorbing (leaking) boundary of tunable strength and a fully reflective boundary. We observe that the absorbing boundary significantly alters the walker’s evolution, with the magnitude of this effect increasing with the absorption strength.
The employed finite system represents an effective minimal model for mimicking energy transport in complex networks, such as those found in biological systems [8, 29]. Here an interesting example is represented by the well-known Fenna-Matthews-Olson (FMO) complex of green sulfur bacteria which is characterized by 7 sites and essentially acts as a molecular wire, transferring excitation energy while showing long-lived quantum coherence [18]. Indeed, FMO complex has been proposed as a dedicated computational device [18], as excitons are able to explore many states simultaneously and efficiently select the correct answer. Our study expands what has been already demonstrated employing the QW approach to introduce further complexity in this dynamics. Precisely, we introduce possible effects due to sites coupling with the environment resulting into energy absorption and decoherence. On the other hand, with this manuscript we explore a field with multiple possible applications because controllable decoherence permits photonic implementations of quantum-computational methods that take advantage of decoherence [24, 25, 47]. In this context, the optimization of suitable transport processes —whether coherent or noise-assisted— may require the engineering of dedicated quantum systems.
@@ 2.1 Theoretical background and simulations
where and represent the probabilities for the walker at site to move at the following step to the left or to the right, respectively, with the normalization condition . The conditional displacement is realized by the shift operator: , which displaces the walker in a superposition of the position basis states according to its internal coin state. As a result, the state of the system at each time step depends recursively on the previous step, i.e. , yielding after steps the coherent evolution , where denotes the initial state of the walker. Throughout this work, we adopted the Hadamard coin operator, , which mimics a fair coin toss. Experimentally, such a coined DTQW can be realized on photonic platforms by injecting photons through a cascade of directional couplers acting as balanced beam splitters (BSs), arranged in a lattice of elementary cells forming Mach–Zehnder (MZ) interferometers. Each BS simultaneously implements both the coin and shift operators, since it splits the photon into left- and right- propagation path and shifts it accordingly. In this scheme, each BS output represents a point in the space-time evolution of the QW: at a given step, the walker is described by the state , where labels the position of the walker along the line corresponding to the spatial mode and denotes the discrete time (). In this configuration, () defines the number of lattice sites available at each time step, while is the maximum number of MZ units per step. Thus the variable takes discrete values given by , where and is an integer index labeling the injection site, with . Building upon our previous results on confined QWs [33], we extended the analysis by introducing an additional key aspect beyond spatial confinement induced by lattice edges, namely the presence of a leaking boundary. We impose asymmetric boundary conditions: one edge of the lattice enforces hard confinement, while the opposite edge acts as a homogeneous leaking boundary. In the former case, the walker is completely reflected at the edge whenever a step would take it outside the lattice. In the latter case, reflection at the boundary is only partial, resulting in a finite probability for the wavefunction to leak out of the lattice, thereby introducing controlled, mode-dependent particle losses (Figure 1).
@@
In order to model this behavior numerically, we constructed a block-diagonal matrix , whose diagonal blocks represented the BS transformations at each time step, with . For the internal sites not corresponding to modes at the edges (), each block consisted of a balanced BS transformation with reflectivity , in order to implement the Hadamard coin. At the lattice edges, the balanced BS was replaced by an unbalanced transformation with reflectivity . In particular, models a perfectly reflective boundary (lossless edge), whereas and model low- and high-leakage scenarios, respectively. At each time step, the block matrix structure accounted for the alternate pattern of an even and an odd number of BSs over subsequent time steps. Here represented the transformation associated with the evolution operator at the -th time step. Hence, the complete evolution of the QW after steps was obtained as the ordered product: . In a QW, the spatial probability profile of the walker after multiple time steps reflects nontrivial interference effects arising from the coherent superposition of the many possible propagation paths. To capture this behavior, we computed the output single-particle probability distributions across all spatial modes, , which represents the probability of finding the particle at position after steps, irrespectively of its internal coin state. In the presence of a leaking boundary, the dynamics is intrinsic non-unitary: since a fraction of the wavefunction is irreversibly lost due to the interaction with leaking edge, the total probability is not longer conserved throughout the time evolution. Therefore, the probability distributions here reported are referred to the population that remained confined within the lattice at each time step. We analyzed the propagation of single photons injected at different initial positions, both close to or far from the leaking boundary, in order to assess the role of the distance from the leaking edge. As quantitative indicators for characterizing the walker dynamics, we computed the time evolution of mean position and its variance , defined as:
@@ 2.2 Experimental setup
The model described above is implemented in a photonic platform. A coherent attenuated beam is used as the walker, while the lattice with asymmetric boundaries is realized using a reconfigurable photonic processor (QuiX Quantum Alquor20) featuring 20 input and output ports [42, 41]. The processor is a fully programmable multiport interferometer capable of implementing arbitrary linear optical transformations over a space whose dimensionality is set by the number of available modes. The interferometer is constructed as a mesh of 190 Mach–Zehnder interferometers (MZIs), each functioning as a tunable beam splitter [12], thereby enabling independent control of the amplitudes and phases of the output signals. The photonic chip is fabricated using stoichiometric silicon nitride () waveguides based on TripleX technology and is mounted on a water-cooled Peltier element to ensure thermal stability. Optical coupling to and from the chip is provided via FC/PC fiber connectors. The processor exhibits an insertion loss of . Its performance is characterized by an average amplitude fidelity of , evaluated over 100 Haar-random unitary matrices at a wavelength of . The fidelity is defined as F = , where denotes the element-wise absolute value and is the number of ports [51]. The fully packaged chip is mounted on a sub-mount to ensure mechanical stability and is electrically interfaced with a printed circuit board (PCB). Control of the device is provided through a Python-based software interface, which allows the user to specify the target transformation and accordingly program the phase shifts applied to the MZIs.
@@ 3.1 Simulations
Due to the confinement, the walker propagation across the lattice follows oscillatory patterns, resulting from a sequence of total or partial reflections at the edges of the accessible region. This behavior is captured by the time evolution of both the mean position and its variance (Figure 3). In the low-loss regime (), the dynamics appears to be only weakly affected by the presence of the leaking boundary and, as shown in Figure 3 (a), the mean position depends on the injection site: when the walker enters the lattice from sites closer to one of the boundaries (input sites 2 and 7, solid lines), it propagates towards the opposite edge, whereas for more internal injection sites (3 and 6, broken lines) the mean position exhibits oscillations of smaller amplitude around the origin of the position axis. Nevertheless, the symmetry associated with the spatial coordinates of the input sites is largely preserved: opposite injection sites (2 and 7, or 3 and 6) share essentially the same temporal evolution, but with opposite phase. Likewise, the time evolution of the variance of the walker’s mean position exhibits an oscillatory behavior (Figure 3 (b)). For each pair of symmetric input sites, the variance follows the same qualitative temporal evolution, indicating that the spreading and partial refocusing of the wavefunction are similarly affected by the lattice boundaries when the leaking probability is low. In particular, when the walker is injected close to either the leaking or the reflecting edge (solid lines), the initial spreading remains limited and the variance stays below 3 for approximately the first 25 steps, before gradually increasing at longer times while preserving fluctuations of large amplitude and high frequency. By contrast, for more internal injection sites the variance displays pronounced oscillations for the earliest steps, corresponding to alternative phases of spreading and relocalization of the wavefunction, occurring with a characteristic period of about 15 steps.
@@
By contrast, in the high-loss regime (), the dynamics of both the mean position and its variance, although still retaining oscillatory features, departs from the regularities and phase relations identified in the low-loss case (Figure 3 (b),(d)). In particular, symmetric injection sites no longer display identical dynamics, and both the amplitude and frequency of the oscillations are modified, indicating a stronger impact of losses on the walker propagation. To further clarify the role of increasing losses on the system dynamics, it is helpful to compare the time evolution of the walker’s mean position and variance for different leaking probabilities and injection sites. The full set of long-time numerical results is presented in Appendix A.
@@
As expected, the strongest sensitivity to losses is observed for input site 2, which is the closest to the leaking boundary (Figure 8 (a)-(b)). As long as the losses remain weak (, blue and green lines), the dynamics closely follow that of the lossless case (, dashed line), with few differences limited to their relative amplitudes up to approximately 40-50 steps. However, when the losses become strong (, red and orange lines), deviations appear already at early times, with a faster propagation of the mean position —for the mean position reaches the first oscillation peak within the first 18 steps, whereas for it requires only 12 steps to reach the first maximum mean displacement— and an enhanced spreading of the wavefunction, as evidenced by the larger variance. For injection sites progressively farther from the leaking boundary (input sites 3, 6, and 7), the impact of losses on the walker’s dynamics appears to be reduced (Figure 8 (c)-(h)). In these configurations, the mean position and variance exhibit similar trends across different leaking regimes for a large number of time steps, with only minor differences in amplitude and phase. This indicates that, when the walker is injected sufficiently far from the lossy edge, the dynamics remain largely governed by coherent propagation within the lattice. In the following, we focus in particular on the initial time steps, for which a direct comparison with experimental observations was feasible.

### v2

@@ Abstract
Quantum walks (QWs) represent pillars of quantum dynamics and information processing. They provide a powerful framework for simulating quantum transport, designing search algorithms, and enabling universal quantum computation. Several physical platforms have been employed for their implementation, such as trapped atoms and ions, nuclear magnetic resonance systems, and photonic quantum architectures either in bulk optics or waveguide structures and fiber-loop networks. Here we focus on the most promising and versatile approach, that is photonic integrated circuits.
In this work, we review how the employment of this versatile experimental platform has allowed to explore several phenomena related to QW-based protocols as, for instance, the evolution in presence of different kinds of noise. In this landscape, to the best of our knowledge, few examples report on the introduction of absorbing centers and their effects on the coherence of the dynamics. Here we present and discuss the results related to absorbing boundaries in QWs obtained through theoretical simulations and experiments conducted with the universal photonic quantum processors realized by Quix Quantum. We analyze how localized absorption along one lattice edge affects the walker dynamics depending on both the leakage probability and the initial injection site. Our results show that the presence of controlled losses modifies interference patterns and coherence, without fully destroying quantum features and providing an effective resource for engineering on-chip QWs and simulating open quantum systems.
@@ 1 Introduction
In real photonic quantum systems, particle losses are unavoidable due to the non-ideal efficiencies of the network components and measurements apparatuses, often representing a major limitation for scaling up quantum applications. Nonetheless, the inclusion of absorbing sites within a QW remains a scenario that has received relatively little attention and has not yet been fully explored. In this regard, a limited number of numerical studies [7, 40, 50, 27, 4] and a few experimental implementations [30, 31] have been reported, showing how some recursive behavior is observed in the presence of specific absorbing sites. However, many aspects of this problem remain largely unexplored, even though the introduction of controlled losses appears to be a promising approach for probing the dynamics of open quantum systems.
In this work, we address this gap by presenting both numerical simulations and experimental results for a QW in the presence of an absorbing boundary. Specifically, we investigate a walk on a finite lattice featuring a partially absorbing (leaking) boundary of tunable strength and a fully reflective boundary. We observe that the absorbing boundary significantly alters the walker’s evolution, with the magnitude of this effect increasing with the absorption strength.
The employed finite system represents an effective minimal model for mimicking energy transport in complex networks, such as those found in biological systems [8, 29]. Here an interesting example is represented by the well-known Fenna-Matthews-Olson (FMO) complex of green sulfur bacteria which is characterized by 7 sites and essentially acts as a molecular wire, transferring excitation energy while showing long-lived quantum coherence [18]. Indeed, the FMO complex has been proposed as a dedicated computational device [18], as excitons are able to explore many states simultaneously and efficiently select the correct answer. Our study expands what has been already demonstrated employing the QW approach to introduce further complexity in this dynamics. Precisely, we introduce possible effects due to sites coupling with the environment resulting into energy absorption and decoherence. On the other hand, with this manuscript we explore a field with multiple possible applications, as controllable decoherence permits photonic implementations of quantum-computational methods that take advantage of decoherence [24, 25, 47]. In this context, the optimization of suitable transport processes —whether coherent or noise-assisted— may require the engineering of dedicated quantum systems.
@@ 2.1 Theoretical background and simulations
where and represent the probabilities for the walker at site to move at the following step to the left or to the right, respectively, with the normalization condition for the coin state . The conditional displacement is realized by the shift operator:
which displaces the walker in a superposition of the position basis states according to its internal coin state. As a result, the state of the system at each time step depends recursively on the previous step, i.e. , yielding after steps the coherent evolution , where denotes the initial state of the walker. Throughout this work, we adopted the Hadamard coin operator, , which mimics a fair coin toss.
@@
Experimentally, such a coined DTQW can be realized on photonic platforms by injecting photons through a cascade of directional couplers acting as balanced beam splitters (BSs), arranged in a lattice of elementary cells forming Mach–Zehnder interferometers (MZIs). Each BS simultaneously implements both the coin and shift operators, since it splits the photon into left- and right- propagation path and shifts it accordingly. In this scheme, each BS output represents a point in the space-time evolution of the QW: at a given step, the walker is described by the state , where labels the position of the walker along the line corresponding to the spatial mode and denotes the discrete time (). In this configuration, () defines the number of lattice sites available at each time step, while is the maximum number of MZI units per step. Thus the variable takes discrete values given by , where and is an integer index labeling the injection site, with .
Building upon our previous results on confined QWs [33], we extended the analysis by introducing an additional key aspect beyond spatial confinement induced by lattice edges, namely the presence of a leaking boundary. We impose asymmetric boundary conditions: one edge of the lattice enforces hard confinement, while the opposite edge acts as a homogeneous leaking boundary. In the former case, the walker is completely reflected at the edge and redirected back into the interior whenever a step would take it outside the lattice. In the latter case, reflection at the boundary is only partial, resulting in a finite probability for the wavefunction to leak out of the lattice, thereby introducing controlled, mode-dependent particle losses (Figure 1).
In order to model this behavior numerically, we constructed a block-diagonal matrix , whose diagonal blocks are represented by the BS transformation
at each time step, with . For the internal sites not corresponding to modes at the edges (), each block given by Equation 3 consisted of a balanced BS transformation with transmissivity , in order to implement the Hadamard coin. At the lattice edges, the balanced BS was replaced by an unbalanced transformation with transmissivity . In particular, models a perfectly reflective boundary (lossless edge), whereas and model low- and high-leakage scenarios, respectively. At each time step, the block matrix structure accounted for the alternate pattern of an even and an odd number of BSs over subsequent time steps. Here represented the transformation associated with the evolution operator at the -th time step. Hence, the complete evolution of the QW after steps was obtained as the ordered product: .
@@
In a QW, the spatial probability profile of the walker after multiple time steps reflects nontrivial interference effects arising from the coherent superposition of the many possible propagation paths. To capture this behavior, we computed the output single-particle probability distributions across all spatial modes, , which represents the probability of finding the particle at position after steps, irrespectively of its internal coin state. In the presence of a leaking boundary, the dynamics is intrinsically non-unitary: since a fraction of the wavefunction is irreversibly lost due to the interaction with leaking edge, the total probability is no longer conserved throughout the time evolution. Therefore, the probability distributions here reported are referred to the population that remained confined within the lattice at each time step. We analyzed the propagation of single photons injected at different initial positions, both close to or far from the leaking boundary, in order to assess the role of the distance from the leaking edge for different leaking probabilities. As quantitative indicators for characterizing the walker dynamics, we computed the time evolution of mean position and its variance , defined as:
@@ 2.2 Experimental setup
The model described above is implemented in a photonic platform. A coherent attenuated beam is used as the walker, while the lattice with asymmetric boundaries is realized using a reconfigurable photonic processor (QuiX Quantum Alquor20) featuring 20 input and output ports [42, 41]. The processor is a fully programmable multiport interferometer capable of implementing arbitrary linear optical transformations over a space whose dimensionality is set by the number of available modes. The interferometer is constructed as a mesh of 190 MZIs interferometers, each functioning as a tunable beam splitter [12], thereby enabling independent control of the amplitudes and phases of the output signals. The photonic chip is fabricated using stoichiometric silicon nitride () waveguides based on TripleX technology and is mounted on a water-cooled Peltier element to ensure thermal stability. Optical coupling to and from the chip is provided via FC/PC fiber connectors. The processor exhibits an insertion loss of . Its performance is characterized by an average amplitude fidelity of , evaluated over 100 Haar-random unitary matrices at a wavelength of . The fidelity is defined as , where denotes the element-wise absolute value and is the number of ports [51]. The fully packaged chip is mounted on a sub-mount to ensure mechanical stability and is electrically interfaced with a printed circuit board (PCB). Control of the device is provided through a Python-based software interface, which allows the user to specify the target transformation and accordingly program the phase shifts applied to the MZIs.
@@ 3.1 Simulations
Due to the confinement, the walker propagation across the lattice follows oscillatory patterns, resulting from a sequence of total or partial reflections at the edges of the accessible region. This behavior is captured by the time evolution of both the mean position and its variance (Figure 4 (a)-(d)). In the low-loss regime (), the dynamics appears to be only weakly affected by the presence of the leaking boundary and, as shown in Figure 4 (a), the mean position depends on the injection site: when the walker enters the lattice from sites closer to one of the boundaries (input sites 2 and 7, cyan dashed and green solid lines), it propagates towards the opposite edge, whereas for more internal injection sites (3 and 6, orange dash-dotted and long-dashed pink lines) the mean position exhibits oscillations of smaller amplitude around the origin of the position axis. Nevertheless, the symmetry associated with the spatial coordinates of the input sites is largely preserved: opposite injection sites (2 and 7, or 3 and 6) share essentially the same temporal evolution, but with opposite phase. Likewise, the time evolution of the variance of the walker’s mean position exhibits an oscillatory behavior (Figure 4 (b)). For each pair of symmetric input sites, the variance follows the same qualitative temporal evolution, indicating that the spreading and partial refocusing of the wavefunction are similarly affected by the lattice boundaries when the leaking probability is low. In particular, when the walker is injected close to either the leaking or the reflecting edge (cyan dashed and green solid lines), the initial spreading remains limited and the variance stays below 3 for approximately the first 25 steps, before gradually increasing at longer times while preserving fluctuations of large amplitude and high frequency. On the other hand, for more internal injection sites the variance displays pronounced oscillations for the earliest steps, corresponding to alternative phases of spreading and relocalization of the wavefunction, occurring with a characteristic period of about 15 steps (orange dash-dotted and long-dashed pink lines).
@@
By contrast, in the high-loss regime (), the dynamics of both the mean position and its variance, although still retaining oscillatory features, departs from the regularities and phase relations identified in the low-loss case (Figure 4 (c),(d)). In particular, symmetric injection sites no longer display identical dynamics, and both the amplitude and frequency of the oscillations are modified, indicating a stronger impact of losses on the walker propagation. To further clarify the role of increasing losses on the system dynamics, it is helpful to compare the time evolution of the walker’s mean position and variance for different leaking probabilities and injection sites. The full set of long-time numerical results is presented in Appendix 5.
@@
As expected, the strongest sensitivity to losses is observed for input site 2, which is the closest to the leaking boundary (Figure 9 (a)-(b)). As long as the losses remain weak (, cyan solid line), the dynamics closely follow that of the lossless case (, gray dotted line), with few differences limited to their relative amplitudes up to approximately 40-50 steps. However, when the losses become strong (, blue dashed line), deviations appear already at early times, with a faster propagation of the mean position (for the mean position reaches the first oscillation peak within the first 18 steps, whereas for it requires only 12 steps to reach the first maximum mean displacement) and an enhanced spreading of the wavefunction, as evidenced by the larger variance. For injection sites progressively farther from the leaking boundary (input sites 3, 6, and 7), the impact of losses on the walker’s dynamics appears to be reduced (Figure 9 (c)-(h)). In these configurations, the mean position and variance exhibit similar trends across different leaking regimes for a large number of time steps, with only minor differences in amplitude and phase. This indicates that, when the walker is injected sufficiently far from the lossy edge, the dynamics remain largely governed by coherent propagation within the lattice. In the following, we focus in particular on the initial time steps, for which a direct comparison with experimental observations was feasible.In order to ensure that boundary effects significantly affected the dynamics, we considered a DTQW on a number of sites small compared to the characteristic diffusion scale of the walker. Specifically, we simulated a QW on a lattice of sites over up to discrete time steps, thereby assessing the long-time behavior. The choice of sites was also guided by the need to allow direct comparison between simulation results and experimental data. We considered four distinct cases depending on the position of the site at which the walker is injected into the lattice: positions close to the leaking boundary (input waveguides 2 and 3, corresponding to the initial states and , respectively), and positions close to the reflective boundary (input waveguides 6 and 7, corresponding to the initial states and , respectively). In our simulations, we adopted for the low-leakage regime and for the high-leakage case. In the present model, particle losses at the leaking boundary are assumed to be homogeneous and time-independent, i.e., the leakage probability is the same at each time step. In Figure 4, the walker mean position and its variance, evaluated over 100 steps, are compared for different input positions and for both the low-loss and high-loss regimes. A more detailed comparison between weak- and strong-leakage dynamics for each input site, including lossless and fully absorbing regimes as reference cases, is reported in Appendix A.
Due to the confinement, the walker propagation across the lattice follows oscillatory patterns, resulting from a sequence of total or partial reflections at the edges of the accessible region. This behavior is captured by the time evolution of both the mean position and its variance (Figure 4 (a)-(d)). In the low-loss regime (), the dynamics appears to be only weakly affected by the presence of the leaking boundary and, as shown in Figure 4 (a), the mean position depends on the injection site: when the walker enters the lattice from sites closer to one of the boundaries (input sites 2 and 7, cyan dashed and green solid lines), it propagates towards the opposite edge, whereas for more internal injection sites (3 and 6, orange dash-dotted and long-dashed pink lines) the mean position exhibits oscillations of smaller amplitude around the origin of the position axis. Nevertheless, the symmetry associated with the spatial coordinates of the input sites is largely preserved: opposite injection sites (2 and 7, or 3 and 6) share essentially the same temporal evolution, but with opposite phase. Likewise, the time evolution of the variance of the walker’s mean position exhibits an oscillatory behavior (Figure 4 (b)). For each pair of symmetric input sites, the variance follows the same qualitative temporal evolution, indicating that the spreading and partial refocusing of the wavefunction are similarly affected by the lattice boundaries when the leaking probability is low. In particular, when the walker is injected close to either the leaking or the reflecting edge (cyan dashed and green solid lines), the initial spreading remains limited and the variance stays below 3 for approximately the first 25 steps, before gradually increasing at longer times while preserving fluctuations of large amplitude and high frequency. On the other hand, for more internal injection sites the variance displays pronounced oscillations for the earliest steps, corresponding to alternative phases of spreading and relocalization of the wavefunction, occurring with a characteristic period of about 15 steps (orange dash-dotted and long-dashed pink lines).
By contrast, in the high-loss regime (), the dynamics of both the mean position and its variance, although still retaining oscillatory features, departs from the regularities and phase relations identified in the low-loss case (Figure 4 (c),(d)). In particular, symmetric injection sites no longer display identical dynamics, and both the amplitude and frequency of the oscillations are modified, indicating a stronger impact of losses on the walker propagation. To further clarify the role of increasing losses on the system dynamics, it is helpful to compare the time evolution of the walker’s mean position and variance for different leaking probabilities and injection sites. The full set of long-time numerical results is presented in Appendix 5.
As expected, the strongest sensitivity to losses is observed for input site 2, which is the closest to the leaking boundary (Figure 9 (a)-(b)). As long as the losses remain weak (, cyan solid line), the dynamics closely follow that of the lossless case (, gray dotted line), with few differences limited to their relative amplitudes up to approximately 40-50 steps. However, when the losses become strong (, blue dashed line), deviations appear already at early times, with a faster propagation of the mean position (for the mean position reaches the first oscillation peak within the first 18 steps, whereas for it requires only 12 steps to reach the first maximum mean displacement) and an enhanced spreading of the wavefunction, as evidenced by the larger variance. For injection sites progressively farther from the leaking boundary (input sites 3, 6, and 7), the impact of losses on the walker’s dynamics appears to be reduced (Figure 9 (c)-(h)). In these configurations, the mean position and variance exhibit similar trends across different leaking regimes for a large number of time steps, with only minor differences in amplitude and phase. This indicates that, when the walker is injected sufficiently far from the lossy edge, the dynamics remain largely governed by coherent propagation within the lattice. In the following, we focus in particular on the initial time steps, for which a direct comparison with experimental observations was feasible.